The functions f and g are defined by f(x) =1/(1-3x) and g(x) =log1/3(3x-2)-log3(x) respectively

1. Write down the sets Df (ehe domain of f) and Dg (the domain of g)

2. Solve the inequality f(x) > 2 for x"\isin" Df

3. Solve the inequality f(x) ≥ 2 for x"\isin" Dg

Hint: Use the change of base formula


1
Expert's answer
2022-01-10T14:43:38-0500

1.



"f(x)=\dfrac{1}{1-3x}""1-3x\not=0=>x\not=\dfrac{1}{3}""Df: (-\infin, \dfrac{1}{3})\cup(\dfrac{1}{3}, \infin)""g(x)=\log_{1/3}(3x-2)-\log_{3}(x)""\begin{cases}
 3x-2>0 \\
 x>0 
\end{cases}=>x>\dfrac{2}{3}""Dg: (\dfrac{2}{3}, \infin)"

2.



"f(x)>2, x\in Df""\dfrac{1}{1-3x}>2, x\in(-\infin, \dfrac{1}{3})\cup(\dfrac{1}{3}, \infin)""\dfrac{1}{1-3x}-2>0""\dfrac{1-2+6x}{1-3x}>0""\dfrac{6x-1}{3x-1}<0""\dfrac{1}{6}<x<\dfrac{1}{3}""x\in(\dfrac{1}{6},\dfrac{1}{3})"

3.



"g(x)\geq2, x\in Dg""\log_{1/3}(3x-2)-\log_{3}(x)\geq2, x\in(\dfrac{2}{3}, \infin)""\log_{1/3}(3x-2)=-\log_{3}(3x-2)""2=\log_{3}(9)""-\log_{3}(3x-2)-\log_{3}(x)\geq\log_{3}(9)""\log_{3}(9x(3x-2))\leq0""3>1=>y(x)=\log_{3}(x)\ increases"

Then



"9x(3x-2)\leq1, x\in(\dfrac{2}{3}, \infin)""27x^2-18x-1\leq0"

Let



"27x^2-18x-1=0""D=(-18)^2-4(27)(-1)=432""x=\dfrac{18\pm\sqrt{432}}{2(27)}=\dfrac{3\pm2\sqrt{3}}{9}=\dfrac{1}{3}\pm\dfrac{2\sqrt{3}}{9}""\begin{cases}
 \dfrac{1}{3}-\dfrac{2\sqrt{3}}{9}\leq x\leq \dfrac{1}{3}+\dfrac{2\sqrt{3}}{9} \\
 x>\dfrac{2}{3} 
\end{cases}""=>\dfrac{2}{3}<x\leq\dfrac{1}{3}+\dfrac{2\sqrt{3}}{9}""x\in\bigg(\dfrac{2}{3},\ \dfrac{1}{3}+\dfrac{2\sqrt{3}}{9}\bigg]"